OpenStudy (anonymous):

Multiply and simplify. (4s)^4*(4s)^12

OpenStudy (agreene):

This one is a bit more tricky because of the parenthesis. Start by distributing to break them: \[(4s)^4*(4s)^{16}=4^4s^4*4^{12}s^{12}\] \[=4^{4+12}s^{4+12}=4^{16}s^{16}\] \[=4294967296 s^{16}\]